cmchar ; beginchar(oct , 10u^{#} + serif fit^{#}, asc height^{#}, desc depth^{#}); italcorr .5x height^{#} ∗ slant + min(.5curve^{#} − .85u^{#}, −.1u^{#}); adjust fit(serif fit^{#}, 0); pickup tiny_{nib}; pos_{1}(stem′, 0); pos_{2}(stem, 0); pos_{0′}(stem′, 0); pos_{0}(stem, 0); z_{0l} = z_{0′l}; x_{0′} = x_{1}; x_{0} = x_{2}; lft x_{1l} = hround(2.5u − .5stem′); top y_{1} = h; numeric edge; edge = rt x_{2r}; pickup fine_{nib}; pos_{3}(if hefty: thin join else: hair fi, 180); pos_{4}(vair′, 90); pos_{5}(curve, 0); pos_{6}(vair, −90); rt x_{3l} = ^{1}/_{3}[rt x_{2}, edge]; y_{3} = ^{1}/_{8}[bar height, x height]; x_{4l} = w − .5(w − serif fit) + .5u; top y_{4r} = x height + oo; rt x_{5r} = hround min(w − 1.35u + .5curve, w − .6u); y_{5} = .5x height; x_{6} = x_{3}; bot y_{6r} = −oo; (x, y_{4r}) = whatever[z_{3l}, z_{4l}]; x_{4r} := min(x, .5[x_{5r}, x_{4}]); filldraw stroke_{z3e}{up} … {right}z_{4e} & super arc_{e}(4, 5)

& z_{5e}{down} … {5(x_{6} − x_{5}), y_{6} − y_{5}}z_{6e}; % bowl y_{0} = ypart(((edge, h) - - (edge, 0)) intersectionpoint (z_{3l}{up} … {right}z_{4l})); pickup tiny_{nib}; bot y_{2} = −d; filldraw stroke_{z1e} - - z_{0′e} - - z_{0e} - - z_{2e}; % stem pickup crisp_{nib}; pos_{8}(hair, 0); pos_{7′}(stem, 0); z_{7′} = z_{2}; x_{8l} = x_{7′l}; bot y_{8} = 0; filldraw stroke_{z7′e} - - z_{8e}; % point if serifs: sloped serif_{l}(1, 0′, a, ^{1}/_{3}, jut, serif drop); % upper serif

dish serif (2, 0, b, ^{1}/_{3}, jut, c, ^{1}/_{3}, 0); fi % lower serif penlabels(0, 1, 2, 3, 4, 5, 6, 8); endchar;